OpenStudy (anonymous):

Quick Computing Company produces calculators. They have found that the cost, c(x), of making x calculators is a quadratic function in terms of x. The company also discovered that it costs $61 to produce 3 calculators, $165 to produce 7 calculators, and $385 to produce 12 calculators. Find the total cost of producing 6 calculators.

OpenStudy (anonymous):

@Jack1 @irishboy123 @TheSadGirl

OpenStudy (jack1):

"cost y(x), of making x calculators is a quadratic function in terms of x" y(x) = ax^2 + bx + c solve for A, B, and C dunno about you, but I'd solve this as 3 simultaneous equations 3 calcs ==> cost = $61 EQN 1. 9a + 3b + c = 61 7 calcs ==> cost = $165 EQN 2. 49a + 7b + c = 165 12 calcs ==> cost = $385 EQN 3. 144a + 12b + c = 385